Sample Questions Relating to Job-Order Costing

  1.     Assume that at the beginning of the year a firm estimates its overhead costs at $400,000. The firm

plans to use job order costing and will use direct labor hours as its base. Direct labor hours for the

year were estimated at 200,000. At the end of the year the firm was under-allocated by $20,000

based on 208,000 actual direct labor hours. What is the actual total overhead cost?

  1.     Cox Company uses a job-order cost system and assigns overhead cost to jobs on a basis of direct

labor hours. Annual manufacturing overhead costs and direct labor hours were estimated at

$120,000 and 40,000 hours, respectively. During the year, Job #334 was completed at a cost of

$10,000 in direct materials and $12,000 in direct labor. The labor rate is $20 per hour. By the end

of the year, Cox Company reported actual work of 38,000 direct labor hours and had incurred

$109,000 in actual manufacturing overhead cost.

  1.     Calculate the total per unit cost of Job #334 which contained 50 units.
  2.     Calculate Cox’s under or overapplied overhead by the end of the year.
  3.     Describe what is causing the under or overapplied overhead.
  1.     Assume that a company estimates the following at the beginning of the year:

Total manufacturing overhead

$1,100,000

Total direct labor cost

$1,120,000

Total direct labor hours

100,000

Total machine hours

50,000

The firm used direct labor hours to apply overhead. If actual overhead at the end of the year is

$1,050,000 and if overhead is overapplied by $215,000, how many direct labor hours were

worked?  
Answer #1

Predetermined rate = Estimated overheads/Estimated Labor Hours

= 400,000/200,000

= $2 per hour

Applied = Actual hours*Predetermined rate

= 208000*2

= $416,000

Actual overheads = Applied + Under allocated

= 416000+20,000

= $436,000

Predetermined Rate = 120,000/40,000 = $3 per hour

Calculation of unit cost

Direct material

10000

Direct labor

12000

Overhead applied

1800

Total cost

23800

Number of units

50

Unit cost

476

Overapplied = 38000*3 - 109000 = $5,000

C. Lower overheads during the year compared to budgeted

Rate per hour = 1,100,000/100,000 = $11 per hour

Applied overhead = 1050000+215000 = $1,265,000

Labor hours worked = 1,265,000/11 = 115000